OpenStudy (anonymous):

3/x-4+2/x+4=14/(x+4)(x-4)

OpenStudy (anonymous):

3/x = (2 (7 x-25))/(x+4)

OpenStudy (anonymous):

Is it to b simplified or what??

OpenStudy (anonymous):

Is it ?? 3 2 14 ---- + ----- = ----------- x-4 x+4 (x+4) (x-4)

OpenStudy (anonymous):

i think you meant: 3/(x-4) +2/(x+4) =14/[(x+4)(x-4)] if yes - need common denominator on left, so: 3*(x+4) +2*(x-4) =14 solve for x... x=2

OpenStudy (anonymous):

x=2